# Zoom-IQA: Image Quality Assessment with Reliable Region-Aware Reasoning

### v1

@@ 10.1 Total Rewards
where $\alpha$ and $\beta$ are coefficients that balance the relative importance of score prediction and rank consistency. In our experiments, we set $\alpha=1$ and $\beta=2$, while the parameter $\sigma$ in the score reward (Eq. 4 in the main paper) is set to 0.35.
@@ 10.2 Reinforcement Learning Training of Self-Guided Exploration
In the Zoom-IQA framework, the VLM first processes the query and the original image. If the initial visual information is insufficient, the model outputs a ”crop” action directed at a specific region, which initiates a subsequent reasoning turn. A localized crop is then fed back into the VLM to generate a refined response. This design establishes an iterative interaction pipeline, enabling the model to progressively resolve ambiguities and refine its assessment. To optimize this multi-turn trajectory, we employ the GRPO [shao2024deepseekmath] objective:
@@
Figure 7: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]) and the baseline (LLaVA-1.6-13b [liu2023visual]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 8: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]) and the baseline (LLaVA-1.6-13b [liu2023visual]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 9: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]) and the baseline (LLaVA-1.6-13b [liu2023visual]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 10: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [li2025q] and VisualQuality-R1 [wu2025visualquality]). We highlight correct descriptions and incorrect descriptions, in addition to the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the blurriness. While both Q-Insight and our method correctly identify both the blurriness and noise, VisualQuality-R1 only recognizes the blurriness and ignores the noise. Crucially, Q-Insight provides only general distortion information, whereas our method not only specifies the distortion types but also precisely identifies the object or region suffering from the distortion. Please zoom in for more details.
@@
Figure 11: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [li2025q] and VisualQuality-R1 [wu2025visualquality]). We highlight correct descriptions and incorrect descriptions, along with the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the digital noise. While Zoom-IQA accurately identifies the issues, VisualQuality-R1 fails to detect minor focus problems and digital noise, claiming there is no significant noise and that details are distinguishable. Furthermore, although Q-Insight notices the softness problem, it reaches a contrastive conclusion (asserting the image has “reasonable detail” but “not a great amount of fine detail to point out”) and also fails to recognize the digital noise. Please zoom in for more details.
@@
Figure 12: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]). We highlight: correct descriptions and incorrect descriptions. The image exhibits minor digital artifacts, which was uniquely identified by Zoom-IQA. Please zoom in for more details.
@@
Figure 13: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]). We highlight: correct descriptions and incorrect descriptions. The image exhibits slight overexposure, which was uniquely identified by Zoom-IQA. Furthermore, VisualQuality-R1 incorrectly classified motion blur as out-of-focus blur. Please zoom in for more details.

### v2

@@ 10.1 Total Rewards
where $\alpha$ and $\beta$ are coefficients that balance the relative importance of score prediction and rank consistency. In our experiments, we set $\alpha=1$ and $\beta=2$, while the parameter $\sigma$ in the score reward (Eq. 4 in main paper) is set to 0.35.
@@ 10.2 Reinforcement Learning Training of Self-Guided Exploration
In the Zoom-IQA framework, the VLM first processes the query and the original image. If the initial visual information is insufficient, the model outputs a ”crop” action directed at a specific region, which initiates a subsequent reasoning turn. A localized crop is then fed back into the VLM to generate a refined response. This design establishes an iterative interaction pipeline, enabling the model to progressively resolve ambiguities and refine its assessment. To optimize this multi-turn trajectory, we employ the GRPO [49] objective:
@@
Figure 7: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 8: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 9: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
@@
Figure 10: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29] and VisualQuality-R1 [67]). We highlight correct descriptions and incorrect descriptions, in addition to the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the blurriness. While both Q-Insight and our method correctly identify both the blurriness and noise, VisualQuality-R1 only recognizes the blurriness and ignores the noise. Crucially, Q-Insight provides only general distortion information, whereas our method not only specifies the distortion types but also precisely identifies the object or region suffering from the distortion. Please zoom in for more details.
@@
Figure 11: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29] and VisualQuality-R1 [67]). We highlight correct descriptions and incorrect descriptions, along with the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the digital noise. While Zoom-IQA accurately identifies the issues, VisualQuality-R1 fails to detect minor focus problems and digital noise, claiming there is no significant noise and that details are distinguishable. Furthermore, although Q-Insight notices the softness problem, it reaches a contrastive conclusion (asserting the image has “reasonable detail” but “not a great amount of fine detail to point out”) and also fails to recognize the digital noise. Please zoom in for more details.
@@
Figure 12: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions and incorrect descriptions. The image exhibits minor digital artifacts, which was uniquely identified by Zoom-IQA. Please zoom in for more details.
@@
Figure 13: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions and incorrect descriptions. The image exhibits slight overexposure, which was uniquely identified by Zoom-IQA. Furthermore, VisualQuality-R1 incorrectly classified motion blur as out-of-focus blur. Please zoom in for more details.